OpenStudy (anonymous):

Prove that the function g(x,y) = [(sin(x^2+y^2)/(sqrt(x^2+y^2)) if (x,y) does not = (0,0) ], [0 if (x,y)=(0,0)] is or is not continuous using the definition of continuity.

OpenStudy (jamesj):

First what's your intuition tell you?

OpenStudy (anonymous):

Well because of the piecewise specifically including 0 and the fact that 0,0 is not a solution I assume that its continuous. But i'm having trouble using the delta epsilon definition of continuity to prove it.

OpenStudy (jamesj):

What do you know about this function f(x) = sin x / x, x not = 0 = 1 , x = 0 Is it continuous?

OpenStudy (anonymous):

yes because the limit as sin(x)/x goes to zero is 1

OpenStudy (anonymous):

Right?

OpenStudy (jamesj):

Yes. Now what about if instead of sin(x)/x, it were sin(x^2)/x ?

OpenStudy (jamesj):

It has limit zero because \[ \frac{\sin x^2}{x} = \frac{\sin x^2}{x^2} x = \left(\frac{ \sin x}{x}\right)^2 x \longrightarrow 0\] as x --> 0 So yes, I think your intuition that the function you have also goes to zero as (x,y) --> 0 and therefore is continuous there. Now to prove it. What kind of proofs are you using? Epsilon-delta?

OpenStudy (jamesj):

sorry, that last bracketed expression is bogus, but even so it still goes to zero as \[ \lim_{x \rightarrow 0} \frac{\sin x^2}{x^2} = 1 \]

OpenStudy (jamesj):

...hence \[ \lim_{x \rightarrow 0} \frac{\sin x^2}{x^2} x = 1 \cdot 0 = 0 \]

OpenStudy (anonymous):

Im using epsilon delta proofs. I still don't quite understand them but I believe that I need to somehow get delta in terms of epsilon

OpenStudy (jamesj):

Correct. Now notice that for x,y sufficiently small, \[ | \sin(x^2 + y^2) | \leq x^2 + y^2 \] That should give you the hint you need.

OpenStudy (anonymous):

So the specific definition that i'm using states If S is a metric space with metric d, ˆ S is a metric space with metric ˆ d, and f : S → ˆ S, then f is continuous at p means that (i) p ∈ S, and (ii) if ǫ is a positive number, then there is a positive number δ such that if q is in S and d(q, p) < δ, then ˆ d(f(q), f(p)) < ǫ. so d(hat) (f(x,y),f(0,0))=If(x,y)-f(0,0)l=I(sin(x^2-y^2)/(sqrt(x^2+y^2))-0I ... following this point an example I have getting it to the form d(a,b)<epsilon but I don't know how to get there with this problem...

OpenStudy (jamesj):

Well, notice that \[ d(0,f(x,y)) = \left| \frac{\sin(x^2 + y^2)}{\sqrt{x^2 + y^2}} \right| \leq \frac{x^2 + y^2}{\sqrt{x^2 + y^2}} = \sqrt{x^2 + y^2} = d(0,(x,y)) \] Now see what to do?

OpenStudy (anonymous):

I see where all the pieces are coming from i just cant seem to straighten it out in my head. so we have d(0,f(x,y)) and that needs to be <epsilon and d(0,(x,y)) which needs to be < delta so sqrt(x^2+y^2) needs to be less than delta and the function needs to be less than epsilon...

OpenStudy (jamesj):

What you want to do is this: given any epsilon > 0, you want to find a delta (which is a function of epsilon) such that d(0,(x,y)) < delta => d(0,f(x,y)) < epsilon We have just shown above something incredibly useful. We have shown that d(0,f(x,y)) < d(0,(x,y)) Hence given any such epsilon, there is an incredibly easy choice of delta such that d(0,(x,y)) < delta => d(0,f(x,y)) < epsilon

OpenStudy (anonymous):

so can we say delta=epsilon?

OpenStudy (jamesj):

Yes. Given any epsilon > 0, choose delta = epsilon. Then d((0,0),(x,y)) < delta = epsilon => d(0,f(x,y)) < epsilon because \[ d(0,f(x,y)) \leq d((0,0),(x,y)) \]

OpenStudy (anonymous):

So that proves that the function is continuous everywhere because you can have a delta for any choice of epsilon?

OpenStudy (jamesj):

No. It proves only that f is continuous at (0,0). But the function is clearly continuous everywhere else, as it is the ratio of continuous functions, the denominator of which is not zero.

OpenStudy (anonymous):

Oh, ok I forgot we were specifically using the point (0,0). Thanks a lot, that really helped to walk through it. We are just given the definition and not really taught how to use it.

OpenStudy (jamesj):

ok. I'm sure in your textbook there are some worked examples and counter-examples. Go and read them carefully, even writing a couple of them out. Epsilon-delta proofs are very subtle and it takes a while to get used to them.

OpenStudy (anonymous):

The text has them in a different format, that is what I tried first. Anyways ill keep practicing.

OpenStudy (jamesj):

And don't feel bad about not "getting it" straight away. It's probably the most subtle kind of proof you've yet seen and lots of people take a while to get used to it. Good luck.

OpenStudy (anonymous):

Thanks!